OpenStudy (thefluffmuncher):

3. Solve, using the substitution method. 3x + 2y = 11 x – 2 = – 4y

OpenStudy (thefluffmuncher):

need help solving this @dtan5457

OpenStudy (anonymous):

Do you want anyone's help? or just dtan5457?

OpenStudy (dtan5457):

3x+2y=11 x-2=-4y Substitution method is to isolate a variable, so that x=, or y=, then to plug into the other equation. x-2=-4y x=-4y+2 Plug that back into first equation 3(-4y+2)+2y=11 -12y+6+2y=11 solve for y -10y=5 y=-0.5 Now you have your y coordinate. Plug y=-0.5 into any of the 2 original equations to get x. Your answer should be (x,-0.5) Just solve for x.